alyssa4453:

Martina's penny bank is 1/6 full. After she adds 440 pennies, it is 2/3 full. How many pennies can Martina's bank hold?

Hero:

Let x represent a full bank. Then \(\dfrac{1}{6}x + 440 = \dfrac{2}{3}x\). Solve for \(x\)